# Sample then Identify: A General Framework for Risk Control and Assessment in Multimodal Large Language Models

### v2

@@ 5 Conclusion
Deploying MLLMs in practical VideoQA applications requires reliable risk control and assessment. In this paper, we present a general two-step framework for risk control by calibrating a conformal score to sample new generations and a nonconformity score to identify high-quality responses, which provides statistically and theoretically rigorous guarantees on the correctness miscoverage rate. Our method bridges the gap between SCP and closed-source MLLMs, and it can also be applied to other generative models (e.g., LLMs) that support sampling in any language generation tasks. Furthermore, we conduct risk assessments and explore the issue of semantic redundancy in prediction sets within open-ended contexts for the first time, which results in a promising evaluation metric for MLLMs in open-ended VideoQA tasks via the average set size after deduplication. Additionally, we analyze reliability measurement in the development of the nonconformity score for more efficient predictions. We hope that our framework can be helpful for human-in-the-loop decision-making and human-AI teams.
@@
Appendix A Background of Conformal Prediction
Conformal prediction (CP) is model-agnostic and distribution-free, and can offer statistically rigorous guarantees of ground-truth coverage on fresh test samples with minimal assumptions (e.g., data exchangeability) given a relatively small calibration set [Angelopoulos and Bates, 2021, Campos et al., 2024]. CP can transform any heuristic notion of uncertainty from any model into a statistically rigorous one by constructing a prediction set adaptively to each particular input.
@@ C.1 Semantic Clustering Process
As noted in Section 3.2 and Algorithm 1, $A_{1}\Leftrightarrow A_{2}$ represents that $A_{1}$ is semantically equivalent to $A_{2}$. Following prior work [Farquhar et al., 2024], we utilize a Natural Language Inference (NLI) classifier for this evaluation. Specifically, we employ the off-the-shelf DeBERTa-large-mnli model [He et al., 2020] as the backbone, and the NLI classifier typically takes $\left[A_{1},A_{2}\right]$ as input and predicts logits for three distinct semantic relationships: entailment, neutral, and contradiction. Note that this function has directionality and we only consider $A_{1}$ and $A_{2}$ to be equivalent when the logit scores corresponding to the entailment are at their highest for both inputs $\left[Q\cup A_{1},Q\cup A_{2}\right]$ and $\left[Q\cup A_{2},Q\cup A_{1}\right]$ (i.e., bidirectional). Here, $Q$ denotes the question, and $\cup$ means we evaluate the semantic relationship between query-response pairs.
@@ Appendix D Datasets
Video-MME (VMME) [Fu et al., 2024] is a multi-modal evaluation benchmark for MLLMs in video analysis, which includes 900 manually selected videos totaling 254 hours and resulting in 2,700 MCQA pairs. NExT-QA (NEXT) [Xiao et al., 2021] is a benchmark for VideoQA that focuses on causal and temporal action reasoning, moving beyond simple scene descriptions. It features both multiple-choice and open-ended QA tasks. However, we only selected the multiple-choice part for our experiments. We use the sub-test set of MCQA, which consists of 2,312 QA pairs. MUSIC-AVQA (MUSC) [Li et al., 2022b] is an open-ended VideoQA dataset designed to evaluate comprehensive multimodal understanding and spatio-temporal reasoning over audio-visual scenes. We use the test set of MUSIC-ACQA, which consists of 9,185 QA pairs. MSVD [Chen and Dolan, 2011] is an open-ended dataset that provides a large-scale, highly parallel text collection. We use 2,519 QA pairs from this dataset. In our experiments, we extracted the audio from the videos and integrated it with the corresponding visual and textual data before inputting them into the MLLMs.
@@ Appendix F Evaluation Metrics
$\mathrm{EER}=\frac{1}{|{\mathcal{D}}_{test}|}\sum_{(x_{i},y_{i})\in{\mathcal{D}}_{test}}\textbf{1}\left\{y_{i}\in{\mathcal{S}}\left(x_{i}\right)\right\}$ (20)
@@
$\mathrm{EER}=\frac{1}{|{\mathcal{D}}_{test}|}\sum_{(x_{i},y_{i})\in{\mathcal{D}}_{test}}\textbf{1}\left\{y_{i}\in{\mathcal{C}}(x_{i})\right\}$ (21)
@@
$\mathrm{ACC}=\frac{1}{|{\mathcal{D}}_{test}|}\sum_{(x_{i},y_{i})\in{\mathcal{D}}_{test}}\textbf{1}\left\{\hat{y}_{ref}^{(i)}=y_{i}\right\}$ (23)
@@ Appendix H Generalizability of TRON
As mentioned before, TRON applies to any generative language models that support sampling (or repeatedly responding to the same query) on various open-ended tasks. To evaluate its generalizability, we employ TRON on the (1) open-book conversational QA dataset, CoQA [Reddy et al., 2019], and the (2) closed-book reading comprehension dataset, TriviaQA [Joshi et al., 2017], utilizing LLMs, LLaMA-3-70B-Instruct [AI@Meta, 2024] and LLaMA-3.1-70B-Instruct. Note that we use LangChain with DeepInfra for LLaMA-3-70B-Instruct  [^5] https://deepinfra.com/meta-llama/Meta-Llama-3-70B-Instruct and LLaMA-3.1-70B-Instruct  [^6] https://deepinfra.com/meta-llama/Meta-Llama-3.1-70B-Instruct. We randomly select 2,000 questions from the development split of CoQA and 2,000 questions from the training split of TriviaQA, and set the split ratio between the calibration and test set to 0.5 for both datasets. Empirical results of the correctness miscoverage rate, obtained from 5 trials with randomly allocated calibration and test data at various user-specified risk levels are shown in Figure 8.
@@
We also utilize the PandaGPT-13B model and GPT-4o  [^7] https://platform.openai.com/docs/models/gpt-4o on the Visual Question Answering (VQA) dataset [Antol et al., 2015] for image understanding tasks. We randomly select 1,200 VQA samples and employ 200 samples as calibration data and 1,000 samples as test data. We set $\alpha$ to 0.1 and evaluate the ERR metric at various user-specified risk levels, $\beta$, as shown in Figure 9.
@@
Empirical results demonstrate that TRON can also provide risk management for both MLLMs and LLMs on image understanding, conversational QA, and reading comprehension tasks. Note that while the theoretical guarantee of correctness coverage is rigorous, there can be minor fluctuations of empirical miscoverage rates that exceed the user-specified risk levels in practice due to finite-sample variability [Angelopoulos and Bates, 2021, Ye et al., 2024].
@@
We also compare TRON with the existing method for LLMs on MCQA tasks, termed Least Ambiguous set-valued Classifiers (LAC[Sadinle et al., 2019, Ye et al., 2024], which has been proven to produce prediction sets with the smallest average size [Sadinle et al., 2019]. LAC defines the nonconformal score function as one minus the softmax score corresponding to the true label. We adopt MMLU [Hendrycks et al., 2020] as the evaluation dataset utilizing LLaMA-3-8B-Instruct  [^8] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct and LLaMA-3.1-8B-Instruct  [^9] https://huggingface.co/meta-llama/Meta-Llama-3.1-8B-Instruct. To maintain consistency with previous work settings [Ye et al., 2024], we only select samples that cover at least one correct response within the candidate set when $M$ is set to 5 (i.e., $\alpha=0$). We set the splitting ratio between calibration and test set to 0.5 and $\beta$ to 0.1, and evaluate the APSS metric. Table 3 demonstrates that when sampling 10 times utilizing the LLaMA-3-8B-Instruct model, TRON’s performance can already match that of LAC, and when $M$ is 20, TRON is more predictive efficient than LAC. Additionally, when the sampling size is set to 10, TRON outperforms LAC. Our insight is that since model internal logits can induce miscalibrated biases due to arbitrarily overfit or otherwise untrustworthy issues, we analyze the output distribution of the model through sampling, which is more sound.

### v3

@@ 5 Conclusion
Deploying MLLMs in practical VideoQA applications requires reliable risk control and assessment. In this paper, we present a general two-step framework for risk control by calibrating a conformal score to sample new generations and a nonconformity score to identify high-quality responses, which provides statistically and theoretically rigorous guarantees on the correctness miscoverage rate. Our method bridges the gap between SCP and closed-source MLLMs, and it can also be applied to other generative models (e.g., LLMs) that support sampling in various language generation tasks (see Appendix H). Furthermore, we conduct risk assessments and explore the issue of semantic redundancy in prediction sets within open-ended contexts for the first time, which results in a promising evaluation metric for MLLMs in open-ended VideoQA tasks via the average set size after deduplication. Additionally, we analyze reliability measurement in the development of the nonconformity score for more efficient predictions. We hope that our framework can be helpful for human-in-the-loop decision-making and human-AI teams.
@@
Acknowledgements
This work was supported by the National Natural Science Foundation of China (Grant No. 62122035) and the project of the Sichuan Provincial Department of Science and Technology (No. 2023YFQ0011). This research was completed during a visiting period at Southern University of Science and Technology.
Appendix A Background of Split Conformal Prediction
Split conformal prediction (SCP) is model-agnostic and distribution-free, and can offer statistically rigorous guarantees of ground-truth coverage on fresh test samples with minimal assumptions (e.g., data exchangeability) given a relatively small calibration set (Angelopoulos & Bates, 2021; Campos et al., 2024). SCP can transform any heuristic notion of uncertainty (i.e., the nonconformity score) from any model into a statistically rigorous one by constructing a prediction set adaptively to each particular input. Additionally, the size of each prediction set allows us to gain intuitive insights into the uncertainty of the current decision-making (i.e., the set becomes larger when the model is uncertain or the input is intrinsically hard) Ye et al. (2024); Kostumov et al. (2024).
@@ C.1 Semantic Clustering Process
As noted in Section 3.2 and Algorithm 1, $A_{1}\Leftrightarrow A_{2}$ represents that $A_{1}$ is semantically equivalent to $A_{2}$. Following prior studies (Duan et al., 2024; Lin et al., 2024; Farquhar et al., 2024), we utilize a Natural Language Inference (NLI) classifier, with the off-the-shelf DeBERTa-large-mnli model (He et al., 2021) as the backbone, for this evaluation. The NLI classifier typically takes $\left[A_{1},A_{2}\right]$ as input and predicts logits for three distinct semantic relationships: entailment, neutral, and contradiction. Note that this function has directionality and we only consider $A_{1}$ and $A_{2}$ to be equivalent when the logit scores corresponding to the entailment are at their highest for both inputs $\left[Q\cup A_{1},Q\cup A_{2}\right]$ and $\left[Q\cup A_{2},Q\cup A_{1}\right]$ (i.e., bidirectional). Here, $Q$ denotes the question, and $\cup$ means we evaluate the semantic relationship between query-response pairs.
@@ Appendix D Datasets
Video-MME (VMME) (Fu et al., 2024) is a multi-modal evaluation benchmark for MLLMs in video analysis, which includes 900 manually selected videos totaling 254 hours and resulting in 2,700 MCQA pairs. NExT-QA (NEXT) (Xiao et al., 2021) is a benchmark for VideoQA that focuses on causal and temporal action reasoning, moving beyond simple scene descriptions. It features both multiple-choice and open-ended QA tasks. However, we only selected the multiple-choice part for our experiments. We use the sub-test set of MCQA, which consists of 2,312 QA pairs. MUSIC-AVQA (MUSC) (Li et al., 2022a) is an open-ended VideoQA dataset designed to evaluate comprehensive multimodal understanding and spatio-temporal reasoning over audio-visual scenes. We use the test set of MUSIC-ACQA, which consists of 9,185 QA pairs. MSVD (Chen & Dolan, 2011) is an open-ended dataset that provides a large-scale, highly parallel text collection. We use 2,519 QA pairs from this dataset. In our experiments, we extracted the audio from the videos and integrated it with the corresponding visual and textual data before inputting them into the MLLMs.
@@ Appendix F Evaluation Metrics
$\mathrm{EER}=\frac{1}{|{\mathcal{D}}_{test}|}\sum_{(x_{i},y_{i})\in{\mathcal{D}}_{test}}\mathbbm{1}\left\{y_{i}\in{\mathcal{S}}\left(x_{i}\right)\right\}$ (20)
@@
$\mathrm{EER}=\frac{1}{|{\mathcal{D}}_{test}|}\sum_{(x_{i},y_{i})\in{\mathcal{D}}_{test}}\mathbbm{1}\left\{y_{i}\in{\mathcal{C}}(x_{i})\right\}$ (21)
@@
$\mathrm{ACC}=\frac{1}{|{\mathcal{D}}_{test}|}\sum_{(x_{i},y_{i})\in{\mathcal{D}}_{test}}\mathbbm{1}\left\{\hat{y}_{ref}^{(i)}=y_{i}\right\}$ (23)
@@ Appendix H Generalizability of TRON
As mentioned before, TRON applies to any generative language models that support sampling (or repeatedly responding to the same query) on various open-ended tasks. To evaluate its generalizability, we employ TRON on the (1) open-book conversational QA dataset, CoQA (Reddy et al., 2019), and the (2) closed-book reading comprehension dataset, TriviaQA (Joshi et al., 2017), utilizing LLMs, LLaMA-3-70B-Instruct (AI@Meta, 2024) and LLaMA-3.1-70B-Instruct. Note that we use LangChain with DeepInfra for LLaMA-3-70B-Instruct  [^5] https://deepinfra.com/meta-llama/Meta-Llama-3-70B-Instruct and LLaMA-3.1-70B-Instruct  [^6] https://deepinfra.com/meta-llama/Meta-Llama-3.1-70B-Instruct. We randomly select 2,000 questions from the development split of CoQA and 2,000 questions from the training split of TriviaQA, and set the split ratio between the calibration and test set to 0.5 for both datasets. Empirical results of the correctness miscoverage rate, obtained from 5 trials with randomly allocated calibration and test data at various user-specified risk levels are shown in Figure 8.
@@
We also utilize the PandaGPT-13B model and GPT-4o  [^7] https://platform.openai.com/docs/models/gpt-4o on the Visual Question Answering (VQA) dataset (Antol et al., 2015) for image understanding tasks. We randomly select 1,200 VQA samples and employ 200 samples as calibration data and 1,000 samples as test data. We set $\alpha$ to 0.1 and evaluate the ERR metric at various user-specified risk levels, $\beta$, as shown in Figure 9.
@@
Empirical results demonstrate that TRON can also provide risk management for both MLLMs and LLMs on image understanding, conversational QA, and reading comprehension tasks. Note that while the theoretical guarantee of correctness coverage is rigorous, there can be minor fluctuations of empirical miscoverage rates that exceed the user-specified risk levels in practice due to finite-sample variability (Angelopoulos & Bates, 2021; Ye et al., 2024).
@@
We also compare TRON with the existing method for LLMs on MCQA tasks, termed Least Ambiguous set-valued Classifiers (LAC(Sadinle et al., 2019; Ye et al., 2024), which has been proven to produce prediction sets with the smallest average size (Sadinle et al., 2019). LAC defines the nonconformal score function as one minus the softmax score corresponding to the true label. We adopt MMLU (Hendrycks et al., 2021) as the evaluation dataset utilizing LLaMA-3-8B-Instruct  [^8] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct and LLaMA-3.1-8B-Instruct  [^9] https://huggingface.co/meta-llama/Meta-Llama-3.1-8B-Instruct. To maintain consistency with previous work settings (Ye et al., 2024), we only select samples that cover at least one correct response within the candidate set when $M$ is set to 5 (i.e., $\alpha=0$). We set the splitting ratio between calibration and test set to 0.5 and $\beta$ to 0.1, and evaluate the APSS metric. Table 3 demonstrates that when sampling 10 times utilizing the LLaMA-3-8B-Instruct model, TRON’s performance can already match that of LAC, and when $M$ is 20, TRON is more predictive efficient than LAC. Additionally, when the sampling size is set to 10, TRON outperforms LAC. Our insight is that since model internal logits can induce miscalibrated biases due to arbitrarily overfit or otherwise untrustworthy issues, we analyze the output distribution of the model through sampling, which is more sound.